# references.bib

### v3

@@ Abstract
The Backpropagation algorithm has often been criticised for its lack of biological realism. In an attempt to find a more biologically plausible alternative, the recently introduced Forward-Forward algorithm replaces the forward and backward passes of Backpropagation with two forward passes. In this work, we show that the internal representations obtained by the Forward-Forward algorithm can organise into category-specific ensembles exhibiting high sparsity – composed of a low number of active units. This situation is reminiscent of what has been observed in cortical sensory areas, where neuronal ensembles are suggested to serve as the functional building blocks for perception and action. Interestingly, while this sparse pattern does not typically arise in models trained with standard Backpropagation, it can emerge in networks trained with Backpropagation on the same objective proposed for the Forward-Forward algorithm. These results suggest that the learning procedure proposed by Forward-Forward may be superior to Backpropagation in modelling learning in the cortex, even when a backward pass is used.
@@ 1 Introduction
Deep Learning is a highly effective approach to artificial intelligence, with tremendous implications for science, technology, culture, and society. At its core, there is the Backpropagation (Backprop) algorithm [rumelhart1986learning], which efficiently computes the gradients necessary to optimise the learnable parameters of an artificial neural network. Backprop, however, lacks biological plausibility [stork1989backpropagation] – leading to many attempts to address the issue. The most recent of such, the Forward-Forward algorithm [hinton2022forward], eliminates the need to store neural activities and propagate error derivatives along the network.
@@
In a standard classification context, the application of Forward-Forward requires the designation of positive and negative data. For example, to classify images, one could assign positive (or negative) data to those images having their correct (or incorrect, respectively) class embedded via one-hot encoding at the border (as shown in Figure 1, Panel A). The Forward-Forward algorithm then learns to discriminate between positive and negative data by optimising a goodness function (e.g., the $\ell_{2}$ norm of the activations), akin to contrastive learning [chen2020simple]. Satisfactory results have been observed [hinton2022forward] for classification tasks on Mnist [lecun2010mnist], a well-known benchmark dataset.
This work goes beyond mere performance, investigating the presence of similarities between biological and artificial neuronal ensembles. Our experiments demonstrate sparsity in representations learned using Forward-Forward, and similarities between these and cortical ensembles found in early stages of sensory processing [miller2014visual]. Neurons that form ensembles are highly specialised, and are found to activate selectively when presented with positive data, while only exhibiting minimal activation to negative data or unstructured stimuli. Our results also indicate that in such representations there may exist collective suppression mechanisms similar to those of biological inhibitory neurons [yuste2015neuron].
Finally, though optimising the cross-entropy loss for the same classification task does not appear to produce the sparse ensembles we observe, the phenomenon may not solely be due to the use of the Forward-Forward algorithm. In fact, similar results are obtained by optimising the same goodness function of Forward-Forward, with Backprop instead. This suggests that more focus should be put on the purpose and biological meaning of the loss function rather than the training algorithm [richards2019deep].
@@
Panel A Examples of activation patterns in response to a typical input. Images show the activation value for network units, arranged as a matrix only for the sake of clarity; darker squares represent more active neurons.
@@
Panel C Activation map for neurons in Layer 1 for all images, grouped by class. A blue dot in position $(x,y)$ indicates that neuron $x$ is activated by input $y$; colorscale represents the intensity of such activation. Horizontal bands mark different categories; blue vertical stripes mark active, category-specific neurons. Each input category activates consistently a specific sets of neurons (ensemble).
@@ 2 Related Work
In the section that follows, we summarise key aspects of the Forward-Forward algorithm and the main findings pertaining identification and characterisation of biological neuronal ensembles in the cortex.
@@ 2.1 Forward-Forward
The Forward-Forward algorithm [hinton2022forward] is a recently proposed learning algorithm for artificial neural networks, whose main premise is the ability to overcome the notorious biological implausibility of Backprop [rumelhart1986learning]. In fact, while the effectiveness of Backprop makes it the standard algorithm for training neural networks, it is based on biologically unrealistic assumptions, such as the need to propagate information forwards and backwards through the network [richards2019deep].
@@
Forward-Forward owes its name to the fact that it replaces the backward pass with an additional forward pass. The two forward passes are executed on different data, named positive and negative data. During training, the objective of Forward-Forward is to maximise a so-called goodness function of the neural activations (e.g. the $\ell_{p}$ norm) on positive data and minimise it on negative data. In a simple image classification setting, such as the one we adopt in this paper, one could encode a class label at the border of images, by one-hot encoding it with a white pixel (as shown in Figure 1, Panel A). Then, following the definition from [hinton2022forward], positive data are those for which the encoded label matches the ground truth label, while the opposite holds for negative data.
Layers are trained separately and sequentially, and learn to discriminate between positive and negative data by maximising and minimising their goodness, according to the data presented. Crucially, activations are normalised before being passed to the subsequent layer, to prevent layers from relying on the goodness computed by their predecessors. From the biological point of view, normalisation is known to be a “canonical neural computation” [carandini2012]. At test time, when a new unlabelled sample has to be categorised, many copies of the image are created, each with a different one-hot encoded label. These are then fed into the neural network to obtain a goodness score. Finally, the image gets classified in the category that produced the maximum goodness value.
In the seminal Forward-Forward paper [hinton2022forward], satisfactory classification results are reported on the standard handwritten digit recognition dataset Mnist, with the definition of positive and negative data described above, and using the $\ell_{2}$ or $\ell_{1}$ norm of activations as goodness function.
In a recent theoretical work, it has been analytically shown that under somewhat mild assumptions sparsity emerges in Forward-Forward layers [yang2023theory] as a consequence of optimizing the Forward-Forward loss. While these theoretical results are derived for a single layer, they bolster the reliability of our experimental findings.
@@ 2.2 Neuronal ensembles
In Neuroscience, ensembles are defined as sparse groups of neurons that co-activate during spontaneous activity or in response to sensory stimuli. Remarkably, ensembles – rather than single neurons – have long been suggested to be emergent functional blocks of cortical activity [yuste2015neuron, hebb2005organization, harris2005neural, buzsaki2010neural] and have a prominent role in sensory processing, memory [hopfield1982neural] and behaviour [carrillo2019controlling]. As shown by [miller2014visual], during visual processing the spiking activity in the cortex is dominated by ensembles, whose properties cannot be accounted for by the independent firing activity of neurons in isolation. Additionally, ensembles can be both generated by sensory stimuli (visual, in this case) or by the spontaneous activity of the network. It is also shown that single neurons can participate in more than one ensemble, thus maximising the encoding potential of the network. Interestingly, the same ensembles are evoked both when reacting to stimuli and during spontaneous activity, indicating that they may serve as inherent fundamental components giving rise to visual responses. Findings from [yoshida2020natural] confirm that visual inputs activate sparse ensembles in the primary visual cortex (V1). Moreover, images can be decoded starting from a low number of highly responsive neurons (i.e. a highly sparse ensemble), and adding more does not increase (or even harms) decoding performance. These neurons consistently and reliably encode the same visual inputs across multiple trials. For that reason, partially overlapping receptive fields may play a role in achieving sparse and robust representation of information. Consequently, collecting the activity of highly responsive neurons emerges as an optimal decoding strategy for downstream neurons. The perdurance of ensembles – they have the capacity to endure for weeks [perez2021long] – suggests that they could also potentially serve as a foundation for the long-term representation of perceptual states or memories.
Finally, the idea of neuronal ensembles has inspired the development of computational models [papadimitriou2020brain]. E.g., [lewicki2004sparse] shows that sparse and redundant (hence more robust) representations are optimal for encoding natural images with noisy or unreliable neurons, a finding as also confirmed by [field1994goal] and [olshausen2004sparse].
@@ 3 Methods
In this work, we investigate and compare the representations produced by three models:
  • A classifier in the style of that used by [hinton2022forward], trained with Forward-Forward (FF);
@@
Such different scenarios are further analysed individually in subsection 3.2, subsection 3.3 and subsection 3.4, respectively.
@@ 3.1 Data
The datasets we use to train and test the models described so far are Mnist [lecun2010mnist], FashionMnist [xiao2017fashion], Svhn [netzer2011reading] and Cifar10 [alex2009learning]. Details on these datasets are provided in the Appendix B.
@@ 3.2 Model trained with Forward-Forward (FF)
Our FF model is inspired by the architecture proposed by [hinton2022forward] – and likewise trained according to the Forward-Forward algorithm. It consists of three fully-connected layers, each composed by $1000$ units in the case of Mnist and FashionMnist, and $3072$ units in the case of Svhn and Cifar10. Each linear layer is followed by elementwise ReLU non-linearities. Both during training and inference, the layer-wise $\ell_{2}$ norm is used as the goodness function of choice; correspondingly, $\ell_{2}$ normalisation is performed between subsequent layers.
@@ 3.3 Model trained with Backpropagation on the goodness objective (BP/FF)
The architecture of the FF model, while designed to be optimised using the Forward-Forward algorithm, can be trained seamlessly with Backpropagation on the same goodness maximisation/minimisation objective. Indeed, keeping the definition of positive and negative data introduced for FF, one could simply use Backprop to optimise the goodness-based loss from the Forward-Forward algorithm.
@@ 3.5 Analysis of representations
For each model described, we analyse the internal representation emerging at each layer. We limit our analysis to data belonging to the test set (i.e. not seen during training) and correctly classified by the respective model. Concretely, the representation of a single image is a $n$-dimensional vector composed by the activations (after the ReLU non-linearity) of all the units in the layer. For each layer, we extract a representation matrix $X$ of size $(M,n)$, where $M$ is the total number of test images (correctly classified) and $n$ is the number of neurons in the layer considered.
@@
For each representation vector $x$ we assign a sparsity measure following the notion of sparsity introduced in [hoyer2004non]:
@@
To detect the emergence of category-specific ensembles, within each model and dataset combination, we adopt the following method. The idea is that a neuron should be considered active and part of an ensemble if it activates consistently and selectively when the network receives input data that belongs to that category.
We start by defining a category-specific representation matrix $X_{c}$, of shape $(M_{c},n)$, where $M_{c}$ is the number of correctly classified test images of the given category. Then, we compute the average activation of each hidden unit across all samples: $\overline{x_{j}}=\frac{1}{M_{c}}\sum_{i=1}^{M_{c}}(X_{c})_{ij}$; and the leave-one-out average of the averages $\text{LOO}_{j}=\frac{1}{n-1}\sum_{i\neq j}\overline{x_{j}}$. We then classify a neuron $i$ as active (i.e. part of an ensemble) if $\overline{x_{i}}>\text{LOO}_{i}$. Examples of average activation profiles and of ensembles are reported in Figure 1.
@@
We can also measure the similarity between two ensembles using the Jaccard similarity index (intersection over union): $J(\mathcal{E}^{c},\mathcal{E}^{c^{\prime}})=\frac{\mid\mathcal{E}^{c}\cap\mathcal{E}^{c^{\prime}}\mid}{\mid\mathcal{E}^{c}\cup\mathcal{E}^{c^{\prime}}\mid}$. Examples of shared units are reported in Figure 3.
When the sparsity $S$ of a representation is low, ensembles are typically ill-defined as too many neurons are significantly active simultaneously and the notion of active unit tend to blur. To set a threshold, we will consider values of $S$ below $0.5$ as not-sparse, and in these cases, we do not define ensembles out of the representation.

### v4

# Emergent representations in networks trained with the Forward-Forward algorithm

@@ Abstract
The Backpropagation algorithm has often been criticised for its lack of biological realism. In an attempt to find a more biologically plausible alternative, the recently introduced Forward-Forward algorithm replaces the forward and backward passes of Backpropagation with two forward passes. In this work, we show that the internal representations obtained by the Forward-Forward algorithm can organise into category-specific ensembles exhibiting high sparsity – composed of a low number of active units. This situation is reminiscent of what has been observed in cortical sensory areas, where neuronal ensembles are suggested to serve as the functional building blocks for perception and action. Interestingly, while this sparse pattern does not typically arise in models trained with standard Backpropagation, it can emerge in networks trained with Backpropagation on the same objective proposed for the Forward-Forward algorithm. footnotetext: Equal contribution. Code is available at https://github.com/NiccoloTosato/EmergentRepresentations
@@ 1 Introduction
Deep Learning is a highly effective approach to artificial intelligence, with tremendous implications for science, technology, culture, and society. At its core, there is the Backpropagation (Backprop) algorithm (Rumelhart et al., 1986), which efficiently computes the gradients necessary to optimise the learnable parameters of an artificial neural network. Backprop, however, lacks biological plausibility (Stork, 1989) – leading to many attempts to address the issue. One of the most recent approaches, the Forward-Forward algorithm (Hinton, 2022), eliminates the need to store neural activities and propagate error derivatives along the network.
@@
In a standard classification context, the application of Forward-Forward requires the designation of positive and negative data. For example, to classify images, one could assign positive (or negative) data to those images having their correct (or incorrect, respectively) class label embedded via one-hot encoding at the border (as shown in Figure 1, Panel A). The Forward-Forward algorithm then learns to discriminate between positive and negative data by optimising a goodness function (e.g., the $\ell_{2}$ norm of the activations), akin to contrastive learning (Chen et al., 2020). Satisfactory results have been observed (Hinton, 2022) for classification tasks on Mnist (Lecun et al., 1998), a standard benchmark dataset. This work takes a step beyond performance evaluation, delving into the structure of the hidden representations learned by the Forward-Forward algorithm, uncovering their spontaneously sparse nature and drawing parallels to neural ensembles observed in the brain (Miller et al., 2014; Yuste et al., 2024).
We organise this paper as follows. In section 2 we set the stage by providing a brief overview of the Forward-Forward algorithm and of neuronal ensembles. Then, section 3 is dedicated to the description of the models and datasets investigated and the methods used to analyse representations. Our analysis of the Forward-Forward representations begins in section 4, where we present our key findings. Specifically, in subsection 4.2, we show that the Forward-Forward algorithm spontaneously learns sparse representations, organised into artificial ensembles, i.e., small sets of highly specialised neurons that consistently co-activate for data in a given class. In subsection 4.3, we demonstrate that these ensembles can overlap, with individual units contributing to multiple ensembles when visual features are shared. Further, subsection 4.4 reveals that ensembles can arise on previously unseen categories, indicating a robust generalization of this representational mechanism. Notably, these ensembles can share units with those associated with seen categories, demonstrating effective integration of new information with concepts learned during training. Finally, in subsection 4.5, we examine the structure of the weights and show that the observed sparsity and ensemble formation arise from suppression mechanisms, analogous to the inhibitory processes mediated by biological neurons (Yuste, 2015). These findings are particularly striking because the Forward-Forward algorithm achieves these properties without requiring explicit regularisation to induce sparsity. We observe that, although optimising the cross-entropy loss for the same classification task does not appear to produce the sparse ensembles we observe, the phenomenon may not solely be due to the use of the Forward-Forward algorithm. In fact, similar results are obtained by optimising the same goodness function of Forward-Forward, with Backprop instead. This suggests that more focus should be put on the purpose and biological meaning of the loss function rather than the training algorithm (Richards et al., 2019). We discuss our results in section 5.
In summary, our main results are as follows:
  • The Forward-Forward algorithm yields sparse representations composed of small groups of highly specific units, which we refer to as ensembles by analogy with those observed in the cortex.
  • Ensembles can emerge in a zero-shot manner for classes held out during training and they can share units across visually related categories.
  • The emergence of sparsity and the formation of ensembles are not unique to Forward-Forward optimisation, as they can also be observed in networks trained with Backpropagation on the same objective function.
@@
Panel A Examples of activation patterns in response to a positive input (class label embedded as a one-hot encoding on the top left corner of the image). Images show the activation value for network units, arranged as a matrix only for the sake of clarity; darker squares represent more active neurons.
@@
Panel C Activation map for neurons in Layer 1 for all images, grouped by class. A blue dot in position $(x,y)$ indicates that neuron $x$ is activated by input $y$; colour scale represents the intensity of such activation. Horizontal bands mark different categories; blue vertical stripes mark active, category-specific neurons. Each input category activates consistently a specific sets of neurons (ensemble).
@@ 2 Related Work
In the section that follows, we summarise key aspects of the Forward-Forward algorithm and the main findings pertaining identification and characterisation of biological neuronal ensembles in the brain.
@@ 2.1 Forward-Forward
The Forward-Forward algorithm (Hinton, 2022) is a recently proposed learning algorithm for artificial neural networks, whose main premise is the ability to overcome the notorious biological implausibility of Backprop (Rumelhart et al., 1986). In fact, while the effectiveness of Backprop makes it the standard algorithm for training neural networks, it is based on biologically unrealistic assumptions, such as the need to propagate information forwards and backwards through the network (Richards et al., 2019).
@@
Forward-Forward owes its name to the fact that it replaces the backward pass with an additional forward pass. The two forward passes are executed on different data, named positive and negative data. During training, the objective of Forward-Forward is to maximise a so-called goodness function of the neural activations (e.g., the $\ell_{p}$ norm) on positive data and minimise it on negative data. In a simple image classification setting, such as the one we adopt in this paper, one could encode a class label at the border of images, by one-hot encoding it with a white pixel (as shown in Figure 1, Panel A). Then, following the definition from Hinton (2022), positive data are those for which the encoded label matches the ground truth label, while the opposite holds for negative data. Layers are trained separately and sequentially, and learn to discriminate between positive and negative data by maximising and minimising their goodness, according to the data presented. Crucially, activations are normalised before being passed to the subsequent layer, to prevent layers from relying on the goodness computed by their predecessors. At test time, when a new unlabelled sample has to be categorised, many copies of the image are created, each with a different one-hot encoded label. These are then fed into the neural network to obtain a goodness score. Finally, the image gets classified in the category that produced the maximum goodness value.
In the seminal Forward-Forward paper (Hinton, 2022), satisfactory classification results are reported on the standard handwritten digit recognition dataset Mnist, with the definition of positive and negative data described above, and using the $\ell_{2}$ or $\ell_{1}$ norm of activations as goodness function. In a recent theoretical work, it has been analytically shown that under somewhat mild assumptions sparsity emerges in Forward-Forward layers (Yang, 2023) as a consequence of optimising the Forward-Forward loss. While these formal results are derived for a single layer, they offer a theoretical grounding for our experimental findings. Recent studies inspired by the Forward-Forward learning procedure have expanded its applicability across various architectures, notably achieving enhanced performance in Convolutional Neural Networks (CNNs) (Papachristodoulou et al., 2024; Sun et al., 2025). Additionally, other works have proposed alternative goodness functions and explored the specific contributions of individual neurons to the classification process, shedding light on the interpretability and adaptability of the approach (Terres-Escudero et al., 2024).
We illustrate properties of representations obtained in Forward-Forward networks, that are reminiscent of what is found the neocortex and hippocampus, where ensembles of a few number of units activate consistently in response to similar stimuli. We discuss properties of neuronal ensembles in the following section.
@@ 2.2 Neuronal ensembles
In Neuroscience, neuronal ensembles are defined as sparse groups of neurons that co-activate either spontaneously or in response to sensory stimuli. These ensembles, rather than individual neurons, have long been proposed as emergent functional units of cortical activity, playing critical roles in sensory processing, memory, and behaviour (Miller et al., 2014; Hebb, 2005; Harris, 2005; György, 2010; Hopfield, 1982; Carrillo-Reid et al., 2019; Carrillo-Reid & Yuste, 2020; Yuste et al., 2024). Recent reviews, such as Yuste et al. (2024), provide a comprehensive overview of the concept and its implications.
The importance of ensembles has been increasingly corroborated by experimental studies, enabled by advances in techniques like calcium imaging, which allow for simultaneous recording of large-scale neural activity at single-cell resolution (Carrillo-Reid & Calderon, 2022). For example, Miller et al. (2014) demonstrated that, during visual processing, cortical spiking activity is dominated by ensembles whose properties cannot be explained by the independent activity of individual neurons. These ensembles are activated both by sensory stimuli (e.g., visual inputs) and by spontaneous network activity, suggesting that they represent intrinsic functional building blocks of cortical responses. Notably, single neurons often participate in multiple ensembles, thereby enhancing the network’s encoding potential (Rigotti et al., 2013; Fusi et al., 2016). Further evidence from Yoshida & Ohki (2020) showed that sparse ensembles in the primary visual cortex (V1) are elicited by visual stimuli. Images can be decoded reliably from the activity of a small subset of highly responsive neurons, with additional neurons either failing to improve or even degrading decoding performance. These findings underscore the efficiency of sparse representations, likely facilitated by partially overlapping receptive fields. This arrangement enables robust and efficient encoding of visual information, making sparse ensembles an optimal strategy for downstream processing.
The presence and functionality of ensembles are not limited to specific species or sensory modalities. Studies in various animal models have revealed their role in diverse neural processes (Dupre & Yuste, 2017; Liu & Baraban, 2019), and recent findings suggest they may even contribute to conscious experience (Boyce et al., 2023). Moreover, ensembles have demonstrated remarkable stability over time. For instance, Pérez-Ortega et al. (2021) showed that neuronal ensembles can persist for weeks, supporting their potential involvement in long-term representations of perceptual states or memories.
Technological advancements have also enabled not just the visualization but the direct stimulation of ensembles, allowing to “play the piano” with ensembles of neurons Carrillo-Reid & Yuste (2020). All-optical approaches, such as those described by Packer et al. (2015) and Carrillo-Reid et al. (2019), have shown that repeatedly stimulating specific groups of neurons in V1 can imprint ensembles that remain spontaneously active even after a day. These imprinted ensembles exhibit pattern completion, where activating a subset of neurons can recall the entire ensemble. Remarkably, this effect persists long after the initial stimulation, and experiments have demonstrated causal links between ensemble activation and behaviour (Carrillo-Reid & Yuste, 2020).
Finally, the concept of neuronal ensembles has inspired computational models. For instance, Doi & Lewicki (2004) demonstrated that sparse and redundant representations are optimal for encoding natural images, particularly when neurons are unreliable, a result corroborated by earlier studies (Field, 1994; Olshausen & Field, 2004). These computational frameworks align with biological observations, suggesting that sparse ensemble representations are both efficient and robust mechanisms for encoding sensory information.
@@ 3 Methods
In this work, we investigate and compare the representations produced by three models  [^1] From this point on, we use the term model to refer to the combination of network architecture and optimisation algorithm.:
  • A classifier in the style of that used by Hinton (2022), trained with Forward-Forward (FF);
@@
Such different scenarios are described individually in subsection 3.2, subsection 3.3 and subsection 3.4, respectively.
@@ 3.1 Data
The datasets we use to train and test the models described so far are Mnist (Lecun et al., 1998), FashionMnist (Xiao et al., 2017), Svhn (Netzer et al., 2011) and Cifar10 (Alex, 2009). Details on these datasets are provided in the Appendix B.
@@ 3.2 Model trained with Forward-Forward (FF)
Our FF model is inspired by the architecture proposed by Hinton (2022) – and likewise trained according to the Forward-Forward algorithm. It consists of three fully-connected layers, each composed by $1000$ units in the case of Mnist and FashionMnist, and $3072$ units in the case of Svhn and Cifar10. Each linear layer is followed by elementwise ReLU non-linearities. Both during training and inference, the layer-wise $\ell_{2}$ norm is used as the goodness function of choice; correspondingly, $\ell_{2}$ normalisation is performed between subsequent layers. Additional results obtained using the $\ell_{1}$ norm as a goodness function are presented in Appendix K.
@@
For comprehensive details on the training procedures of the models discussed in this section and the next two sections, we direct the reader to Appendix C.
@@ 3.3 Model trained with Backpropagation on the goodness objective (BP/FF)
The architecture of the FF model, while designed to be optimised using the Forward-Forward algorithm, can be trained seamlessly with Backprop on the same goodness maximisation/minimisation objective. Indeed, keeping the definition of positive and negative data introduced for FF, one could simply use Backprop to optimise the goodness-based loss from the Forward-Forward algorithm.
@@ 3.5 Analysis of representations
For each model described, we analyse the internal representation emerging at each layer. We limit our analysis to data belonging to the test set (i.e., not seen during training) and correctly classified by the respective model. However, the main results of our analysis, concerning sparse and ensemble-like representations, extend without any modification to training data. Concretely, the representation of a single image is a $n$-dimensional vector composed by the activations (after the ReLU non-linearity) of all the units in the layer. For each layer, we extract a representation matrix $X$ of size $(M,n)$, where $M$ is the total number of test images (correctly classified) and $n$ is the number of neurons in the layer considered.
@@
For each representation vector $x$ we assign a sparsity measure following the notion of sparsity introduced in Hoyer (2004):
@@
To detect the emergence of category-specific ensembles, within each model and dataset combination, we adopt the following method. The idea is that a neuron should be considered active and part of an ensemble if it activates consistently and specifically when the network receives input data that belongs to that category.
We start by defining a category-specific representation matrix $X_{c}$, of shape $(M_{c},n)$, where $M_{c}$ is the number of correctly classified test images of the given category. Then, we compute the average activation of each hidden unit across all samples: $\overline{x_{j,c}}=\frac{1}{M_{c}}\sum_{i=1}^{M_{c}}(X_{c})_{ij}$; and the leave-one-out average of the averages $\text{LOO}_{j,c}=\frac{1}{n-1}\sum_{i\neq j}\overline{x_{j,c}}$. We then classify a neuron $i$ as active (i.e., part of an ensemble) if $\overline{x_{i,c}}>2\cdot\text{LOO}_{i,c}$. We also perform a significance test for these comparisons through a permutation test (see Appendix D and Table 3). Examples of average activation profiles and of ensembles are reported in Figure 1.
@@
We can also measure the similarity between two ensembles using the Jaccard similarity index (intersection over union): $J(\mathcal{E}^{c},\mathcal{E}^{c^{\prime}})=\frac{\mid\mathcal{E}^{c}\cap\mathcal{E}^{c^{\prime}}\mid}{\mid\mathcal{E}^{c}\cup\mathcal{E}^{c^{\prime}}\mid}$. As an example, for two ensembles composed of 50 units with a substantial ensemble overlap of $30\%$ (15 units), the similarity $J$ is $\approx 0.18$. Examples of shared units are reported in Figure 3 (see Table 3 for typical ensemble sizes in our setting).
When the sparsity $S$ of a representation is low, ensembles are typically ill-defined as too many neurons are significantly active simultaneously and the notion of active unit tends to blur. To set a threshold, we will consider values of $S$ below $0.5$ as non-sparse, and in these cases, we do not define ensembles out of the representation.